OpenStudy (anonymous):

. How many grams of water can be prepared from 5 moles of hydrogen at standard conditions? 2 H2 (g) + O2(g) → 2 H2O

OpenStudy (aaronq):

you need to use the coefficients and moles to relate what you want: \(\sf \dfrac{moles~of~hydrogen~gas}{H_2~'s~coefficient}=\dfrac{moles~of~water}{H_2O's~coefficient}\) plug your values then solve. Next convert the moles of water to grams (mass): \(\sf moles=\dfrac{mass}{Molar~mass}\)